The three angles in a triangle always add up to 180°. If one angle in a triangle is 48° and the second is 2 times the third, what are the three angles?
Answer by checkley77(12844)

48+2*48+x=180
48+96+x=180
144+x=180
x=180-144
x=36 degrees ans. for the third angle.